# Localization and splitting of a quantum droplet by immersing a heavy impurity

### v1

@@ IV Droplet dynamics
The analysis of the existence and stability properties of static droplet configurations under the influence of repulsive or attractive central potential well paves the way to probe specific dynamical features of the original harmonically trapped droplet. This investigation is certainly relevant for guiding possible experimental endeavors aiming to explore the effect of the potential well on the droplet distributions. For instance, according to the above-discussed phenomenology it is possible to utilize the repulsive barrier to observe droplet fragmentation or the attractive potential to excite internal modes of the droplet. Also, the BdG excitation spectrum can be used to identify the nature of the dynamically excited modes. Below, we showcase such dynamical traits with the aid of a quench of the potential well on the initial harmonically trapped droplet distribution.
@@ IV.1 Dynamical splitting of the droplet
First, we examine the quench-induced dynamics caused by a sudden ramp-up of the repulsive well characterized by specific widths, $\sigma$, and heights, $V_{0}$. To be concrete, we consider an initially harmonically trapped droplet with $\mu=-0.2$ (i.e. close to the lower $\mu$ limit of existence of droplet solutions) experiencing an external potential described by Eq. (2) with $V_{0}=0$ and $\Omega=0.001$. Subsequently, we quench both $V_{0}>0$ and $\sigma$ to realize a double-well potential as shown in Fig. 1(a). Revisiting the corresponding stationary configurations illustrated in Fig. 2(d)-(f), it is anticipated that such a quench facilitates, in general, the splitting of the droplet into two segments. In what follows, we initially discuss quenches of the same potential height, $V_{0}=1$, and different widths $\sigma$ providing an overview of the emergent dynamical features. Subsequently, generalizations of our arguments are also offered for varying height of the potential barrier.
@@
To appreciate the impact of the potential width on the dynamical response, we present the time-evolution of the droplet’s density for fixed $V_{0}=1$ and different postquench values of $\sigma$ in Figs. 4(a), (d). It can be observed that a relatively small width, e.g. $\sigma=0.05\ll\xi$ shown in Fig. 4(a), is not sufficient to split the original droplet. This is anticipated from the relevant ground-state distributions where for $\sigma=0.05$ the droplet fragments after $V_{0}>4.5$. However, the ramping of the repulsive barrier suffices to significantly perturb the droplet distribution by imprinting a permanent density dip at the droplet core, see also the relevant density profiles at different time instants in Fig. 4(b), since atoms are repelled from this region. Notice here the resemblance between the effect of the repulsive barrier on the atom distribution with the impact of repulsive impurities immersed in a 1D Bose gas Grusdt et al. (2024); Mistakidis et al. (2023). Additionally, the presence of the potential well triggers a collective motion of the droplet. This practically refers to a breathing mode of the droplet whose distribution periodically expands and contracts [see also Fig. 4(b)] possessing a frequency $\omega\approx 0.125\approx\omega_{R}^{(2)}$. We have checked that this value matches the prediction of the suitable branch appearing in the corresponding real part of the appropriate BdG spectrum (not shown).
@@
On the other hand, increasing the potential well’s width results in a dramatically different dynamical response, see Fig. 4(d). Indeed, larger widths facilitate droplet separation due to the increased force exerted by the potential, in line with our predictions at the stationary level [see Fig. 2(f)]. Therefore, the original harmonically trapped droplet gradually splits into two counterpropagating fragments traveling towards the trap edges as shown in Figs. 4(d), (e). These are droplet fragments, as evidenced by the fitting of their waveforms to the analytical droplet solution [Eq. (5)] shown in Fig. 4(e) at $t=120$. Simultaneously, each of the nucleated droplet fragments becomes excited as can be inferred from its vibrational pattern over time in Fig. 4(d). This excitation dynamics of the droplet fragments is also accompanied by the emission of relatively small density portions, see in particular the blue faint density ripples surrounding the droplet segments in the course of the evolution. This phenomenon is a manifestation of the droplet’s self evaporation Ferioli et al. (2020) due to its excited nature.
Figure 6: Time-evolution of the quenched droplet position variance, $\braket{X^{2}(t)}/\braket{X^{2}(0)}$, for (a) fixed height, $V_{0}=-1$ and various widths (see legend) as well as (c) a fixed width, $\sigma=0.5$, and different heights (see legend). The recurring behavior of the variance signals the collective mode excitation of the droplet with decaying amplitude. For relatively large quenches, the system tends to a quasi-steady state at long evolution times captured by the almost constant behavior of the variance for $t>50$. The frequency spectrum of the position variance is depicted in panel (b) [(d)] for distinct widths [heights] (see legends). The well-defined peaks in the spectrum refer to the collective motion of the droplet in the modified harmonic trap. For all settings, $\mu=-0.2$ and $\delta=1$.
@@
In order to track the droplet trajectory or the distance among the nucleated droplets depending on the characteristics of the quenched repulsive potential well we compute
[…]
@@
Here, $L$ denotes the size of our computational domain and the factor of two takes into account the fact that the counterpropagating droplet fragments are symmetric with respect to $x=0$, implying that $D(t)$ measures their distance in time. This observable can be experimentally monitored via in-situ measurements Ronzheimer et al. (2013). The time-evolution of $D(t)$ is shown in Fig. 4(c), (f) for various postquench potential amplitudes, $V_{0}$, and widths, $\sigma$. It can be easily deduced that for quenches of fixed width, e.g. $\sigma=0.5\sim\xi$, but different potential height the distance among the generated droplets increases over time in an almost linear manner, see Fig. 4(c). Particularly, larger amplitude quenches yield in general a larger inter-droplet separation accompanied by an increased velocity of the nucleated droplets. This is related to the fact that potentials with larger heights exert a stronger force on the original droplet resulting in more violent separation. The relatively small amplitude oscillations imprinted on $D(t)$ essentially reflect the excited nature of the nucleated droplets, see also Fig. 4(d). A similar response takes place by assuming fixed height of the repulsive potential and variable width [see Fig. 4(f)]; increasing $\sigma$ leads to larger separations and velocities of the generated droplets. This is again traced back to the increased force from the barrier to the droplet.
[…]
Figure 7: Droplet density dynamics using a quench of the central attractive potential well with height, $V_{0}=-20$, from width $\sigma=0$ to (a) $\sigma=0.05$ and (c) $\sigma=0.5$. Self-evaporation is observed in both cases with the small width quench resulting in a quasi-steady state and the large width one to a highly excited droplet. (b) [(d)] Density profiles of (a) [(c)] at selected time-instants (see legends). Panels (b), (d) showcase fits of the analytical droplet solution [Eq. (5)] to the eGPE distribution at distinct evolution times on the density level (see legend). Good agreement is observed at the droplet core. Here, the droplet parameters are the same as in Fig. 6.
@@ IV.2 Droplet localization and self-evaporation
Next, we proceed by considering quenches of the attractive potential well atop the initially harmonically trapped droplet. This protocol is expected to induce dynamical localization of the droplet structure since it favors the trapping of a certain number of atoms into the bound state created by the central well. Such a behavior is, of course, further supported by the ground state analysis of the system illustrated in Fig. 2(a)-(c).
@@
Characteristic examples of the emergent droplet density dynamics after a quench to fixed $V_{0}=-1$ and distinct potential widths are demonstrated in Fig. 5(a), (c). Focusing on relatively small widths, such as $\sigma=0.05\ll\xi$, we observe a localization tendency of the perturbed droplet where a significant portion of atoms resides in the vicinity of the attractive potential, see Fig. 5(a), performing a breathing motion. The aforementioned localization trend of the droplet can be clearly discerned in the instantaneous density profiles, displayed in Fig. 5(b), exhibiting an overall spatially shrinked distribution. The latter clearly interchanges, during the evolution, between a configuration having a spatially distorted peak structure [see e.g. the profile at $t=0.2$ in panel (b)] around $x=0$ where the attractive potential lies, and a “triangle” shaped density resembling a peakon [see, for instance, panel (b) at $t=100$]. The latter structures [^3] They are also reminiscent of the bound state solution of a single-particle in an attractive delta potential ($\psi(x)\sim\sqrt{\kappa}e^{-\kappa\left|x\right|}$, $\kappa=\left|V_{0}\right|\sigma\sqrt{2\pi})$, see also the discussion in Sec. II. are similar to the analytical droplet solution in the presence of a delta potential as can be inferred from the fits provided in Fig. 5(b). Moreover, the above-discussed response is reminiscent of the reaction of a bosonic host gas attractively interacting with impurity atoms (especially of infinite mass), giving rise to attractive polaron states Grusdt et al. (2024); Mistakidis et al. (2020).
@@
The accumulation of atoms towards the minimum of the attractive potential causes also a collective motion of the droplet cloud within the deformed trap corresponding to a breathing motion with frequency $\omega\approx 0.18$. The latter can be estimated by exploiting the position variance of the bosonic cloud which is given by
$\braket{X^{2}(t)}=\braket{\psi(x,t)}{\hat{x}^{2}}{\psi(x,t)}.$ (8)
[…]
@@
Further increasing the width of the attractive potential, e.g. to $\sigma=0.5\sim\xi$ results in a comparatively more violent initial evolution and a subsequent prominent localization. This can be deduced from the large amplitude of $\braket{X^{2}(t)}/\braket{X^{2}(0)}$ and its gradual decaying oscillatory pattern [characterized by a large frequency band with a small amplitude, Fig. 6(b)] toward a constant value that signals the approach to a quasi-steady state at long times. The emergent complex response can be easily visualized in the density evolution displayed in Fig. 5(c). Shortly after the quench, the majority of atoms are dynamically trapped by the attractive potential forming a central excited droplet [see also Fig. 5(d)]. This central droplet is described to a good approximation (except from the tails) by the analytical solution in free space [Eq. (5)] as inferred by the fitted waveforms illustrated in Fig. 5(d). The remaining atoms escape from the potential dip and form two counterpropagating density branches (matter jets) traveling to the edges, see e.g. the density profile at $t=5$ in Fig. 5(d). Subsequently, the excited droplet undergoes a relaxation process by emitting small density fractions, reaching eventually a quasi-steady state. Such a behavior is reminiscent of the self-evaporation mechanism observed in three-dimensional droplets Ferioli et al. (2020); Fort and Modugno (2021).
Turning to larger values of the postquench potential height, e.g. $V_{0}=-20$, we observe a similar to the above-described (for $V_{0}=-1$, $\sigma=0.5\sim\xi$) droplet response, see Fig. 7(a), (c). More concretely, independently of the width of the potential, an initial localization takes place accompanied by the emission of two counterpropagating density segments that can not be trapped by the attractive potential. It is, however, interesting that at least for short-times the droplet core can be captured to a certain extent by the analytical droplet solution of Eq. (5), see the fits in Fig. 7(b), (d). Nevertheless, the emitted density segments are naturally more prominent for smaller widths, compare the elevated density tails of $\left|\psi\right|^{2}$ in Fig. 7(b) with the ones in panel (d). Following this process, a localized central droplet is produced whose excited nature can be deduced from the undulated instantaneous density profiles depicted in Fig. 7(b), (d). The degree of excitation depends also on the potential widths by means that wider ones imprint a larger amount of excitations which are in fact retained in the course of the evolution, see the vibrating density in Fig. 7(c) as compared to the far less disturbed one in Fig. 7(a).
To further inspect the dynamical relaxation tendency of the quenched droplet with respect to the height of the potential well, we present the time-evolution of the position variance, $\braket{X^{2}(t)}/\braket{X^{2}(0)}$, for different $V_{0}$ and fixed $\sigma$ in Fig. 6(c). It turns out that in all cases there is an initial decaying tendency of the recurring behavior of the droplet cloud at short times which is altered for quite large $V_{0}$ values, such as $V_{0}=-20$. For the latter quench the produced excitation is persistent enforcing internal vibrations of the droplet which performs a breathing motion. The characteristic frequencies of this mode are identified in the spectral peaks of the position variance shown in Fig. 6(d). These peaks clearly shift to larger frequencies for increasing $V_{0}$. According to the above description, we can infer that the characteristics of the attractive potential well can be used not only to control the degree of droplet localization, and hence its atom number, but also its excitation frequency and dictate whether it will approach a quasi-steady state in the long time-evolution.
@@
Figure 8: High amplitude probabilities of finding the droplet in distinct single-particle eigenstates (see legend) of the external potential described by Eq. (2) in the course of the evolution. The potential well used is (a) repulsive with characteristics $V_{0}=1$, $\sigma=0.05$, and attractive with (b) $V_{0}=-1$, $\sigma=0.05$, (c) $V_{0}=-5$, $\sigma=0.05$ and (d) $V_{0}=-20$, $\sigma=0.5$. For a relatively shallow postquench repulsive or attractive potential, several eigenstates are populated whose periodic trend is associated with the collective motion of the droplet [panels (a), (b)]. Larger potential heights lead to the occupation of a reduced number of eigenstates but of higher amplitude reflecting either the tendency to a quasi-steady state [panel (c)] or the high excitation of the droplet [panel (d)]. In all cases, the chemical potential corresponds to $\mu=-0.2$ and the LHY strength is $\delta=1$.

### v2

@@ IV Droplet dynamics
The analysis of the existence and stability properties of static droplet configurations under the influence of a central potential barrier or well paves the way to probe specific dynamical features of the original harmonically trapped droplet. This investigation is certainly relevant for guiding possible experimental endeavors aiming to explore the effect of the potential defect on the droplet distributions. For instance, according to the above-discussed phenomenology it is possible to utilize the potential barrier to observe droplet fragmentation or the potential well to excite internal modes of the droplet. Also, the BdG excitation spectrum can be used to identify the nature of the dynamically excited modes. Below, we showcase such dynamical traits with the aid of a quench of the potential defect on the initial harmonically trapped droplet distribution.
@@ IV.1 Dynamical splitting of the droplet
First, we examine the quench-induced dynamics caused by a sudden ramp-up of the potential barrier characterized by specific widths, $\sigma$, and heights, $V_{0}$. To be concrete, we consider an initially harmonically trapped droplet with $\mu=-0.2$ (i.e. close to the lower $\mu$ limit of existence of droplet solutions) experiencing an external potential described by Eq. (2) with $V_{0}=0$ and $\Omega=0.001$. Subsequently, we quench both $V_{0}>0$ and $\sigma$ to realize a double-well potential as shown in Fig. 1(a). Revisiting the corresponding stationary configurations illustrated in Fig. 2(d)-(f), it is anticipated that such a quench facilitates, in general, the splitting of the droplet into two segments. In what follows, we initially discuss quenches of the same potential height, $V_{0}=1$, and different widths $\sigma$ providing an overview of the emergent dynamical features. Subsequently, generalizations of our arguments are also offered for varying height of the potential barrier.
@@
To appreciate the impact of the potential width on the dynamical response, we present the time-evolution of the droplet’s density for fixed $V_{0}=1$ and different postquench values of $\sigma$ in Figs. 4(a), (d). It can be observed that a relatively small width, e.g. $\sigma=0.05\ll\xi$ shown in Fig. 4(a), is not sufficient to split the original droplet. This is anticipated from the relevant ground-state distributions where for $\sigma=0.05$ the droplet fragments after $V_{0}>4.5$. However, the ramping of the potential barrier suffices to significantly perturb the droplet distribution by imprinting a permanent density dip at the droplet core, see also the relevant density profiles at different time instants in Fig. 4(b), since atoms are repelled from this region. Notice here the resemblance between the effect of the potential barrier on the atom distribution with the impact of repulsive impurities immersed in a 1D Bose gas Grusdt et al. (2024); Mistakidis et al. (2023). Additionally, the presence of the potential well triggers a collective motion of the droplet. This practically refers to a breathing mode of the droplet whose distribution periodically expands and contracts [see also Fig. 4(b)] possessing a frequency $\omega\approx 0.125\approx\omega_{R}^{(2)}$. We have checked that this value matches the prediction of the suitable branch appearing in the corresponding real part of the appropriate BdG spectrum (not shown).
@@
On the other hand, increasing the potential barrier’s width results in a dramatically different dynamical response, see Fig. 4(d). Indeed, larger widths facilitate droplet separation due to the increased force exerted by the potential, in line with our predictions at the stationary level [see Fig. 2(f)]. Therefore, the original harmonically trapped droplet gradually splits into two counterpropagating fragments traveling towards the trap edges as shown in Figs. 4(d), (e). These are droplet fragments, as evidenced by the fitting of their waveforms to the analytical droplet solution [Eq. (5)] shown in Fig. 4(e) at $t=120$. Simultaneously, each of the nucleated droplet fragments becomes excited as can be inferred from its vibrational pattern over time in Fig. 4(d). This excitation dynamics of the droplet fragments is also accompanied by the emission of relatively small density portions, see in particular the blue faint density ripples surrounding the droplet segments in the course of the evolution. This phenomenon is a manifestation of the droplet’s self evaporation Ferioli et al. (2020) due to its excited nature.
Figure 5: Dynamics of the droplet density following a quench of the central potential well from height $V_{0}=0$ to $V_{0}=-1$ for various widths (a) $\sigma=0.05$ and (c) $\sigma=0.5$. Relative small amplitude quenches trigger a collective motion of the droplet [panel (a)]. However, larger amplitudes lead to self-evaporation along with droplet excitation at short times which eventually decays tending to a quasi-steady state at long times [panel (c)]. (b) [(d)] Instantaneous densities taken from (a) [(c)]. Panels (b), (d) demonstrate also fits of the analytical droplet solution described by Eq. (4) and Eq. (5) respectively, to the eGPE obtained wave functions at specific time-instants (see legend). Adequate agreement occurs, excluding the tails of the distributions. The chemical potential used is $\mu=-0,2$ and the strength of the LHY $\delta=1$.
@@
In order to track the droplet trajectory or the distance among the nucleated droplets depending on the characteristics of the quenched potential barrier we compute
$D(t)=2\int_{0}^{L/2}dx~x~\left|\psi(x,t)\right|^{2}.$ (7)
@@
Here, $L$ denotes the size of our computational domain and the factor of two takes into account the fact that the counterpropagating droplet fragments are symmetric with respect to $x=0$, implying that $D(t)$ measures their distance in time. This observable can be experimentally monitored via in-situ measurements Ronzheimer et al. (2013). The time-evolution of $D(t)$ is shown in Fig. 4(c), (f) for various postquench potential barrier amplitudes, $V_{0}$, and widths, $\sigma$. It can be easily deduced that for quenches of fixed width, e.g. $\sigma=0.5\sim\xi$, but different potential barrier height the distance among the generated droplets increases over time in an almost linear manner, see Fig. 4(c). Particularly, larger amplitude quenches yield in general a larger inter-droplet separation accompanied by an increased velocity of the nucleated droplets. This is related to the fact that barriers with larger heights exert a stronger force on the original droplet resulting in more violent separation. The relatively small amplitude oscillations imprinted on $D(t)$ essentially reflect the excited nature of the nucleated droplets, see also Fig. 4(d). A similar response takes place by assuming fixed height of the potential barrier and variable width [see Fig. 4(f)]; increasing $\sigma$ leads to larger separations and velocities of the generated droplets. This is again traced back to the increased force from the barrier to the droplet.
An energy argument can be employed to intuitively understand the inter-droplet separation, namely that the potential energy after the barrier quench is converted to kinetic for the two fragments. The potential energy, $E_{p}=\int dx~V(x)$ $\left|\psi(x,0)\right|^{2}$, can be approximated as $E_{p}\simeq\left|\psi(0,0)\right|^{2}V_{0}\sigma\sqrt{2\pi}$, where we assumed that the initial droplet is constant within $\left|x\right|<\sigma$. Subsequently, $E_{p}$ is equally partitioned into the two droplet fragments, each acquiring a velocity $v\simeq\sqrt{E_{p}}$. This implies that $D(t)=2vt$ is a good approximation only for large $\sigma$ and moderate $V_{0}$. For the other cases, the droplet fragments become significantly excited, and thus the above energy argument becomes invalid. Summarizing, we can infer that tuning either the height or the width of the potential barrier provides a knob to control the inter-droplet distance and each fragment’s velocity.
Figure 6: Time-evolution of the quenched droplet position variance, $\braket{X^{2}(t)}/\braket{X^{2}(0)}$, for (a) fixed height, $V_{0}=-1$ and various widths (see legend) as well as (c) a fixed width, $\sigma=0.5$, and different heights (see legend) of the potential well. The recurring behavior of the variance signals the collective mode excitation of the droplet with decaying amplitude. For relatively large quenches, the system tends to a quasi-steady state at long evolution times captured by the almost constant behavior of the variance for $t>50$. The frequency spectrum of the position variance is depicted in panel (b) [(d)] for distinct widths [heights] (see legends). The well-defined peaks in the spectrum refer to the collective motion of the droplet in the modified harmonic trap. For all settings, $\mu=-0.2$ and $\delta=1$.
@@ IV.2 Droplet localization and self-evaporation
Next, we proceed by considering quenches of the potential well atop the initially harmonically trapped droplet. This protocol is expected to induce dynamical localization of the droplet structure since it favors the trapping of a certain number of atoms into the bound state created by the central well. Such a behavior is, of course, further supported by the ground state analysis of the system illustrated in Fig. 2(a)-(c).
Figure 7: Droplet density dynamics using a quench of the central potential well with height, $V_{0}=-20$, from width $\sigma=0$ to (a) $\sigma=0.05$ and (c) $\sigma=0.5$. Self-evaporation is observed in both cases with the small width quench resulting in a quasi-steady state and the large width one to a highly excited droplet. (b) [(d)] Density profiles of (a) [(c)] at selected time-instants (see legends). Panels (b), (d) showcase fits of the analytical droplet solution [Eq. (5)] to the eGPE distribution at distinct evolution times on the density level (see legend). Good agreement is observed at the droplet core. Here, the droplet parameters are the same as in Fig. 6.
@@
Characteristic examples of the emergent droplet density dynamics after a quench to fixed $V_{0}=-1$ and distinct potential well widths are demonstrated in Fig. 5(a), (c). Focusing on relatively small widths, such as $\sigma=0.05\ll\xi$, we observe a localization tendency of the perturbed droplet where a significant portion of atoms resides in the vicinity of the potential well, see Fig. 5(a), performing a breathing motion. The aforementioned localization trend of the droplet can be clearly discerned in the instantaneous density profiles, displayed in Fig. 5(b), exhibiting an overall spatially shrinked distribution. The latter clearly interchanges, during the evolution, between a configuration having a spatially distorted peak structure [see e.g. the profile at $t=0.2$ in panel (b)] around $x=0$ where the potential well lies, and a “triangle” shaped density resembling a peakon [see, for instance, panel (b) at $t=100$]. The latter structures [^3] They are also reminiscent of the bound state solution of a single-particle in an attractive delta potential ($\psi(x)\sim\sqrt{\kappa}e^{-\kappa\left|x\right|}$, $\kappa=\left|V_{0}\right|\sigma\sqrt{2\pi})$, see also the discussion in Sec. II. are similar to the analytical droplet solution in the presence of a delta potential as can be inferred from the fits provided in Fig. 5(b). Moreover, the above-discussed response is reminiscent of the reaction of a bosonic host gas attractively interacting with impurity atoms (especially of infinite mass), giving rise to attractive polaron states Grusdt et al. (2024); Mistakidis et al. (2020).
@@
The accumulation of atoms towards the minimum of the potential well causes also a collective motion of the droplet cloud within the deformed trap corresponding to a breathing motion with frequency $\omega\approx 0.18$. The latter can be estimated by exploiting the position variance of the bosonic cloud which is given by
$\braket{X^{2}(t)}=\braket{\psi(x,t)|\hat{x}^{2}|\psi(x,t)}.$ (8)
In this expression, we assume that $\braket{\psi(x,t)|\hat{x}|\psi(x,t)}=0$ since quenching the potential well preserves the parity symmetry of the droplet cloud with respect to $x=0$, a condition that has been also numerically confirmed. Note also that this observable can be experimentally probed through in-situ absorption measurements Ronzheimer et al. (2013). The time-evolution of $\braket{X^{2}(t)}/\braket{X^{2}(0)}$ for the particular quench is presented in Fig. 6(a). It clearly captures the periodic expansion and contraction dynamics of the droplet cloud. Moreover, its frequency spectrum depicted in Fig. 6(b), $f(\omega)=\int dt~e^{-i\omega t}\braket{X^{2}(t)}/\braket{X^{2}(0)}$, shows a clear peak at $\omega\approx 0.18\approx\omega^{(2)}_{R}$ referring to the breathing mode and being in line with the BdG predictions, see $\omega^{(2)}_{R}$ for $V_{0}=-1$ in Fig. 3(a). This periodic response of the droplet cloud persists when maintaining the same potential height but increasing its postquench width, e.g. to $\sigma=0.1<\xi$. However, the spectrum of $\braket{X^{2}(t)}/\braket{X^{2}(0)}$ exhibits a slightly larger frequency, see Fig. 6(b). An outcome that is supported by our BdG predictions dictating that for fixed $V_{0}$, larger $\sigma$ values “push” the breathing frequency also to larger values.
@@
Further increasing the width of the potential well, e.g. to $\sigma=0.5\sim\xi$ results in a comparatively more violent initial evolution and a subsequent prominent localization. This can be deduced from the large amplitude of $\braket{X^{2}(t)}/\braket{X^{2}(0)}$ and its gradual decaying oscillatory pattern [characterized by a large frequency band with a small amplitude, Fig. 6(b)] towards a constant value that signals the approach to a quasi-steady state at long times. The emergent complex response can be easily visualized in the density evolution displayed in Fig. 5(c). Shortly after the quench, the majority of atoms are dynamically trapped by the potential well forming a central excited droplet [see also Fig. 5(d)]. This central droplet is described to a good approximation (except from the tails) by the analytical solution in free space [Eq. (5)] as inferred by the fitted waveforms illustrated in Fig. 5(d). The remaining atoms escape from the potential well and form two counterpropagating density branches (matter jets) traveling to the edges, see e.g. the density profile at $t=5$ in Fig. 5(d). Subsequently, the excited droplet undergoes a relaxation process by emitting small density fractions, reaching eventually a quasi-steady state. Such a behavior is reminiscent of the self-evaporation mechanism observed in three-dimensional droplets Ferioli et al. (2020); Fort and Modugno (2021).
@@
Figure 8: High amplitude probabilities of finding the droplet in distinct single-particle eigenstates (see legend) of the external potential described by Eq. (2) in the course of the evolution. (a) The potential barrier used has characteristics $V_{0}=1$, $\sigma=0.05$, and the potential well is described by (b) $V_{0}=-1$, $\sigma=0.05$, (c) $V_{0}=-5$, $\sigma=0.05$ and (d) $V_{0}=-20$, $\sigma=0.5$. For a relatively shallow postquench potential barrier or well, several eigenstates are populated whose periodic trend is associated with the collective motion of the droplet [panels (a), (b)]. Larger potential heights lead to the occupation of a reduced number of eigenstates but of higher amplitude reflecting either the tendency to a quasi-steady state [panel (c)] or the high excitation of the droplet [panel (d)]. In all cases, the chemical potential corresponds to $\mu=-0.2$ and the LHY strength is $\delta=1$.
Turning to larger values of the postquench potential well height, e.g. $V_{0}=-20$, we observe a similar to the above-described (for $V_{0}=-1$, $\sigma=0.5\sim\xi$) droplet response, see Fig. 7(a), (c). More concretely, independently of the width of the potential well, an initial localization takes place accompanied by the emission of two counterpropagating density segments that cannot be trapped by the potential well. It is, however, interesting that at least for short-times the droplet core can be captured to a certain extent by the analytical droplet solution of Eq. (5), see the fits in Fig. 7(b), (d). Nevertheless, the emitted density segments are naturally more prominent for smaller widths, compare the elevated density tails of $\left|\psi\right|^{2}$ in Fig. 7(b) with the ones in panel (d). Following this process, a localized central droplet is produced whose excited nature can be deduced from the undulated instantaneous density profiles depicted in Fig. 7(b), (d). The degree of excitation depends also on the potential widths by means that wider ones imprint a larger amount of excitations which are in fact retained in the course of the evolution, see the vibrating density in Fig. 7(c) as compared to the far less disturbed one in Fig. 7(a).
To further inspect the dynamical relaxation tendency of the quenched droplet with respect to the height of the potential well, we present the time-evolution of the position variance, $\braket{X^{2}(t)}/\braket{X^{2}(0)}$, for different $V_{0}$ and fixed $\sigma$ in Fig. 6(c). It turns out that in all cases there is an initial decaying tendency of the recurring behavior of the droplet cloud at short times which is altered for quite large $V_{0}$ values, such as $V_{0}=-20$. For the latter quench the produced excitation is persistent enforcing internal vibrations of the droplet which performs a breathing motion. The characteristic frequencies of this mode are identified in the spectral peaks of the position variance shown in Fig. 6(d). These peaks clearly shift to larger frequencies for increasing $V_{0}$. According to the above description, we can infer that the characteristics of the potential well can be used not only to control the degree of droplet localization, and hence its atom number, but also its excitation frequency and dictate whether it will approach a quasi-steady state in the long time-evolution.